A rectangle has a length of 38 yards less than 9 times its width. If the area of the rectangle is 5992 spare yards. Find the length of the rectangle
Answer by Timnewman(323)

let the width of the rectngle
be w
lenght L=9w-38 ----(1)
Are of the a rectangle=L*W
therefore,
LW=5992 ---(2)
From (1)
L=9w-38
w=(L+38)/9 ----(3)
substitute for w in (2) using(3)
L(L+38)/9=5992
(Lē+38L)/9=5992
Lē+38L=53928
then,
Lē+38L-53892=0
solving the quadratic equation,
L=309.3yerds